W. Oevel Mathematik für Physiker III

Veranstaltungsnr: 172090 Skript zur Vorlesung, Universität Paderborn, Wintersemester 2002/2003

Zeit und Ort: V2 Mi 9.15 – 10.45 D1.312
V2 Fr 11.15 – 12.45 E2.304
Ü2 Mo 9 – 11  E2.304 (Kai Gehrs)

Zur Homepage der Veranstaltung:

math-www.uni-paderborn.de.de/∼walter (→ Lehrveranstaltungen)
Email: walter@upb.de

Inhalt

1 Fourier–Analysis 1 1.1 Motivation: Ein erstes Beispiel . . . . . . . . . . . . . . . . . . . 1 1.2 Approximation im quadratischen Mittel . . . . . . . . . . . . . . 7

1.2.1 Skalarprodukte und Hilbert–Räume . . . . . . . . . . . . 7 1.2.2 Orthogonale Systeme . . . . . . . . . . . . . . . . . . . . . 12 1.2.3 Bestapproximation und Konvergenz im quadratischen

Mittel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

1.3 Fourier–Reihen . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 1.3.1 Definition der trigonometrischen Fourier–Reihen . . . . . 19 1.3.2 Konvergenz im quadratischen Mittel . . . . . . . . . . . . 22 1.3.3 Punktweise Konvergenz . . . . . . . . . . . . . . . . . . . 24 1.3.4 Abklingverhalten der Fourier–Koeffizienten . . . . . . . . 33 1.3.5 Das Gibbs’sche Phänomen . . . . . . . . . . . . . . . . . . 36 1.3.6 Diskrete Fourier–Transformation (DFT) . . . . . . . . . . 40 1.3.7 Schnelle Fourier–Transformation (FFT) . . . . . . . . . . 43 1.3.8 Technische Durchführung der FFT . . . . . . . . . . . . . 46 1.3.9 Beliebige Perioden . . . . . . . . . . . . . . . . . . . . . . 53 1.3.10 Mehrdimensionale Fourier–Transformation . . . . . . . . . 54 1.3.11 Einige Anwendungen . . . . . . . . . . . . . . . . . . . . . 56

1.4 Die Fourier–Integraltransformation . . . . . . . . . . . . . . . . . 63 1.4.1 Motivation . . . . . . . . . . . . . . . . . . . . . . . . . . 63 1.4.2 L_{1}–Theorie . . . . . . . . . . . . . . . . . . . . . . . . . . 64 1.4.3 L_{2}–Theorie . . . . . . . . . . . . . . . . . . . . . . . . . . 72 1.4.4 Distributionen und ihre Fourier–Transformation . . . . . 78 1.4.5 Die Unschärferelation . . . . . . . . . . . . . . . . . . . . 82 1.4.6 Zusammenfassung . . . . . . . . . . . . . . . . . . . . . . 87

2 Sturm–Liouville–Probleme 89 2.1 Das Eigenwertproblem . . . . . . . . . . . . . . . . . . . . . . . . 89 2.2 Reguläre Sturm–Liouville–Probleme . . . . . . . . . . . . . . . . 104 2.3 Inverse Sturm–Liouville–Operatoren . . . . . . . . . . . . . . . . 107

2.4 Anwendungen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112

3 Komplexe Funktionen 117 3.1 Differenzierbarkeit in ℂ . . . . . . . . . . . . . . . . . . . . . . . 117 3.2 Konturintegrale in ℂ . . . . . . . . . . . . . . . . . . . . . . . . . 120 3.3 Der Cauchysche Integralsatz . . . . . . . . . . . . . . . . . . . . . 126 3.4 Potenzreihen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 3.5 Der Residuenkalkül . . . . . . . . . . . . . . . . . . . . . . . . . . 136

4 Laplace–Transformation 145 4.1 Strukturelles: Existenz und Eindeutigkeit . . . . . . . . . . . . . 145 4.2 Rechenregeln . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 4.3 Rücktransformation . . . . . . . . . . . . . . . . . . . . . . . . . 158

4.3.1 Vereinfachungen über Rechenregeln . . . . . . . . . . . . . 158 4.3.2 Systematische Rücktransformation . . . . . . . . . . . . . 161 4.3.3 Numerische Rücktransformation . . . . . . . . . . . . . . 168

4.4 Anwendung: die ”Systemtheorie“ der Ingenieure . . . . . . . . . . 172

Literatur

Die Vorlesung baut nicht streng auf irgendeinem Buch auf, sondern geht ihren eigenen Weg. Die angegebenen Referenzen dienen dazu, sich unabhängig vom Skript entsprechende Grundlagen anzueignen oder spezielle Inhalte zu vertiefen. Es handelt sich um eine recht willkürliche Auswahl: Neben den angegebenen Büchern gibt es sicherlich jede Menge weiterer Literatur, die den behandelten Stoff analog abdeckt.

Zu Kapitel 1 (Fourier–Analysis):

1) H. Dym and H. P. McKean : Fourier Series and Integrals, Academic Press, 1972 This book contains numerous applications of Fourier analysis. Strongly recommended for anyone who is interested in applications and wants to deepen their understanding of Fourier analysis. It also includes a nice description of Lebesgue integration and group theory.

2) T. W. Körner : Fourier Analysis, Cambridge University Press, 1988 This is a monumental work on Fourier analysis, consisting of a bunch of interrelated essays. Read one section per day! You will gain a lot. Highly recommended.

3) J. S. Walker : Fourier Analysis, Oxford University Press, 1988 A well-written and solid book on Fourier analysis with applications on optics, computer-aided tomography, spherical harmonics, etc.

4) G. B. Folland : Fourier Analysis and Its Applications, Brooks/Cole Publishing Co., 1992 An introductory but extremely well-written textbook on Fourier analysis. Contains chapters on special functions, generalized functions (distributions), and Greens functions. Applications are mainly for differential equations. Expensive but worth buying it.

5) J. M. Ash (ed.) : Studies in Harmonic Analysis, Mathematical Association of America, 1976 This is a collection of conference talks by the authorities held in Chicago in 1975. Most of the chapters are as if these authorities are directly talking to you in a friendly manner about the essence of the ideas in harmonic analysis without much detailed proofs. Contains really deep mathematics.

6) S. G. Krantz : A Panorama of Harmonic Analysis, Mathematical Association of America, 1999 This book gives a historical perspective of harmonic analysis ranging from classical to modern, from elementary to advanced. One can see how subtle it is to sum multiple Fourier series. This also includes short description on wavelets. Highly recommended.

7) E. M. Stein and G. Weiss : Introduction to Fourier Analysis on Euclidean Spaces, Princeton University Press, 1971 A classic of the multidimensional Fourier analysis. Includes detailed discussions on the invariance properties of Fourier transform.

8) A. Zygmund : Trigonometric Serie, (2nd Ed., Volume I & II combined), Cambridge University Press, 1959 An ultimate bible on Fourier series and integrals for hard analysts. This is basically a dictionary. Almost no applications are treated here.

9) R. N. Bracewell : The Fourier Transform and Its Application, (2nd Ed., Revised), MacGraw-Hill, 1986 Another bible for engineers. Contains an excellent pictorial dictionary of many functions and their Fourier transforms.

10) G. P. Tolstov : Fourier Series, Dover, 1972.

The most cost effective book (about $12). Very well written. Highly recommended.

11) G. H. Hardy and W. W. Rogosinski : Fourier Series, Dover, 1999. This is a prelude to Zygmund’s book. Spirit of pure mathematics. No applications included. Economical ($7).

12) W. L. Briggs and V. E. Henson : The DFT: An Owner’s Manual for the Discrete Fourier Transform, SIAM 1995 This is a very useful book on DFT. Includes many practical applications, such as tomography, seismic migrations, difference equation solvers. Detailed analysis on the error of the DFT. A nice book to keep on your desk.

13) A. Terras : Fourier Analysis on Finite Groups and Applications, Cambridge University Press, 1999. Another type of Fourier analysis. A more detailed version of the first half of Chapter 4 of Dym and McKean plus many more examples and applications of that aspect of Fourier analysis.

Zu Kapitel 2 (Sturm–Liouville–Probleme):

[WW] Wolfgang Walter, Gewöhnliche Differentialgleichungen, Springer, 1996.

[Tol] G. P. Tolstov : Fourier Series, Dover, 1972. Man beachte hier speziell das letzte Kapitel.

Zu Kapitel 3 (Komplexe Funktionen):

[Jän] Klaus Jänich: Funktionentheorie – Eine Einführung, Springer, 1999.

Zu Kapitel 4 (Laplace–Transformation):

[Föl] Otto Föllinger, Laplace–, Fourier– und z–Transformation, Hüthig Verlag, 2000.

[Mar] Jerrold E. Marsden, Basic Complex Analysis, Freeman 1987.

[BR] R.E. Bellmann and R.S. Roth, The Laplace Transform, World Scientific 1984.

[Doe] Gustav Doetsch, Einführung in Theorie und Anwendung der Laplace– Transformation, Birkhäuser 1970.

[SGV] W. Strampp, V. Ganzha, E. Vorozhtsov, Höhere Mathematik mit Mathematica 4, Vieweg, 1997.